Henson Company applies overhead on the basis of 120% of direct labor cost. Job No. 190 is charged with $80,000 of direct materials costs and $120,000 of manufacturing overhead. The total manufacturing costs for Job No. 190 is

A $200,000
B $344,000
C $216,000
D $300,000
Answer #1
Concepts and reason

Job Costing: This type of costing is used when goods are produced based upon the specifications expected from the customer at different situations. When goods are produced based upon the expectations of the customer is known as Job order costing.

It is essential for determining the cost of the job. In this case, Job costing is used to estimate the cost of the manufacturing the products as required by the customers. Therefore it is clear that cost for each job differ from one another.

Job costing usually involves in keeping a track of direct costs and indirect costs that are related with the particular job. In this method, overheads that are related to the jobs are allocated based on Direct Labor Hours.

Predetermined overhead rate: The cost accountants estimate a rate for all the overhead costs to compute manufacturing overhead. That allocation rate, which is computed as a ratio of total estimated overhead costs to total estimated quantity, is known as predetermined overhead rate.

Fundamentals

Pre-Determined Overhead rate: The predetermined overhead allocation rate is used to allocate the total overhead costs to the direct labor costs.

Total Manufacturing Costs: The costs which do not relate directly with the manufacturing of products are referred to as manufacturing overhead costs or indirect costs.

Formula for Total Manufacturing costs are as follows:

TotalManufacturingCosts=DirectMaterials+DirectLabor+ManufacturingOverhead{\rm{Total}}\,{\rm{Manufacturing}}\,{\rm{Costs = }}\,{\rm{Direct}}\,{\rm{Materials}}\,{\rm{ + }}\,{\rm{Direct}}\,{\rm{Labor}}\,{\rm{ + }}\,{\rm{Manufacturing}}\,{\rm{Overhead}}

Direct materials: Direct materials are the raw materials which are directly related with the production of the goods.

Direct labor cost: the labor cost includes the wages or salaries paid to employees who physically produce goods or services. Labor refers to the actual work that the employees do to produce products. Labor costs refer to the amount of money that the employer pays the employees to complete the work.

Directlaborhourcost=Estimatedoverhead×100%Predeterminedoverheadrate{\rm{Direct labor hour cost = Estimated overhead }} \times {\rm{ }}\frac{{100\% }}{{{\rm{Predetermined overhead rate}}}}

Manufacturing overhead costs: The costs, which do not relate directly with the manufacturing of products, are referred to as manufacturing overhead costs or indirect costs.

Calculate direct labor using predetermined overhead and manufacturing overhead.

Directlaborhourcost=Estimatedoverhead×100%Predeterminedoverheadrate=$120,000×100%120%=$100,000\begin{array}{c}\\{\rm{Direct labor hour cost = Estimated overhead }} \times {\rm{ }}\frac{{100\% }}{{{\rm{Predetermined overhead rate}}}}\\\\ = \,\$ 120,000\, \times \,\frac{{100\,\% }}{{120\,\% }}\,\\\\ = \$ 100,000\\\end{array}

Therefore, direct labor hours cost is $100,000.

Calculate Total Manufacturing Costs as shown below.

H COMΡΑNY
TOTAL MANUFACTURING COSTS FOR JOB NO. 190
Particulars
Amount
Direct materials
Direct labor
Manufacturing overhead
T

Ans:

Total manufacturing costs for Job No. 190is $300,000.